OpenStudy (chris215):

Given f '(x) = (5 - x)(8 - x), determine the intervals on which f(x) is increasing or decreasing.

zepdrix (zepdrix):

increasing decreasing? So we need to determine where the function "turns" from increasing to decreasing and all that jazz, ya? We call those locations "critical" or "stationary" points.

zepdrix (zepdrix):

critical points exist where the first derivative is zero. \[\large\rm 0=(5-x)(8-x)\]

zepdrix (zepdrix):

So what values do you get for critical points? :) We can construct a first derivative test once you've determined those points.

OpenStudy (chris215):

5 and 8?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 0>(5-x)(8-x) }\) when f decreases (negative slope) \(\color{#000000 }{ \displaystyle 0<(5-x)(8-x) }\) when f increases (positive slope)

OpenStudy (solomonzelman):

solve each thing for x.

OpenStudy (triciaal):